Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
verkindt - 19:31 Wednesday 15 October 2025 (67951) Print this report
Comment to Glitch on the horizon visible in B1_DC_NULL (67950)

Adding a TIME plot showing B1_PD1_DC, B1_PD2_DC, B1_PD1_Audio, B1_PD2_Audio, B1_DC_NULL and LSC_DARM.
The glitch covers a frequency band up to 300 Hz and has a large low frequency component.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×